OpenStudy (anonymous):

a recent survey by a local radio station shows that 3of 8 listeners like a new song while 1/4 hate it. the rest have no strong opinion. a) if 1600 people answered the survey, how many like the new song? dislike the new song? have no opinion? b) could these results apply to the entire listening audience? explain

OpenStudy (anonymous):

1600*3/8=600 1600*1/4=400 1600-600-400=600 b)no not all listeners could be surveyed